# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ Abstract
In this paper, we analyze the sample complexities of learning the optimal state-action value function $Q^{*}$ and an optimal policy $\pi^{*}$ in a finite discounted Markov decision process (MDP) where the agent has recursive entropic risk-preferences with risk-parameter $\beta\neq 0$ and where a generative model of the MDP is available. We provide and analyze a simple model based approach which we call model-based risk-sensitive $Q$-value-iteration (MB-RS-QVI) which leads to $(\varepsilon,\delta)$-PAC-bounds on $\|Q^{*}-Q^{k}\|$, and $\|V^{*}-V^{\pi_{k}}\|$ where $Q_{k}$ is the output of MB-RS-QVI after k iterations and $\pi_{k}$ is the greedy policy with respect to $Q_{k}$. Both PAC-bounds have exponential dependence on the effective horizon $\frac{1}{1-\gamma}$ and the strength of this dependence grows with the learners risk-sensitivity $|\beta|$. We also provide two lower bounds which shows that exponential dependence on $|\beta|\frac{1}{1-\gamma}$ is unavoidable in both cases. The lower bounds reveal that the PAC-bounds are tight in the parameters $S,A,\delta,\varepsilon$ and that unlike in the classical setting it is not possible to have polynomial dependence in all model parameters.
@@ 1 Introduction
In reinforcement learning (RL), the aim of the agent is to conventionally optimize the expected return, which is defined in terms of a (discounted) sum of rewards sutton1998reinforcement . In the majority of RL literature, the environment is modeled via the Markov Decision Process (MDP) framework puterman2014markov , wherein efficient computation of an optimal policy, thanks to optimal Bellman equations, renders possible. However, as a risk-neutral objective, the expected return fails to capture the true needs of many high-stake applications arising in, e.g., medical treatment ernst2006clinical , finance scutella2013robust ; bielecki1999risk , and operations research delage2010percentile . Decision making in such applications must take into account the variability of returns, and risks thereof. To account for this, one may opt to to maximize a risk measure of the return distribution, while another approach could be to consider the entire distribution of return, as is done under the distributional RL framework bellemare_distributional_2023 that has received much attention over the last decade.
@@
Within the first approach, the risk is quantified via concave risk measures, which yield well-defined mathematical optimization frameworks. Notably, they include value-at-risk (VaR), Conditional VaR (CVaR) shapiro2021LNs , entropic risk howard1972risk , and entropic VaR (EVaR) ahmadi2012entropic , all of which have been applied to a wide-range of scenarios. CVaR appears to be the most popular one used to model risk-sensitivity in MDPs chow2014algorithms ; bisi2022risk ; brown2020bayesian ; bauerle_2011_markov , mainly due to a delicate control it offers for the undesirable tail of return. Despite its popularity and rich interpretation, solving and learning MDPs with CVaR-defined objectives has rendered technically challenging bauerle_2011_markov . This has been a key motivation of adopting weaker notions such as nested CVaR bauerle_markov_2022 , at the expense of sacrificing the interpretability. Entropic risk, as another popular notion, has been considered for risk control in MDPs borkar2002risk ; neu2017unified ; hau2023entropic ; hu2023tighter ; fei2020risk . In the RL literature, it has been mainly considered for the undiscounted settings, despite the popularity of discounted MDPs. A notable exception is hau2023entropic that studies, among other things, planning in discounted MDPs under the entropic risk.
In this paper, we study risk-sensitive discounted RL where the agent’s objective is formulated using the entropic risk measure. In discounted RL, where future rewards are discounted by a factor of $\gamma$, one may identify two main approaches to apply the entropic risk to sequence of rewards, $(r_{t})_{t\geq 0}$, collected by an RL agent. The first and most intuitive one, which we may call the non-recursive approach, consists in directly applying the entropic risk functional to return $\sum_{t=0}^{\infty}\gamma^{t}r_{t}$. The other approach, which we may call the recursive approach, works by applying the risk functional at every step $t$ (see Section 2 for details). The non-recursive approach (e.g., hau2023entropic ), while being most intuitive, has several drawbacks; e.g., the optimal policy might not be time-consistent (see jaquette1976utility ). In contrast, the recursive approach yields a form of Bellman optimality equation, which is key to developing learning algorithms with provable sample complexity. Therefore, we restrict attention to the RL problem defined using recursive risk-preferences.
1.1 Main Contributions
@@
We study risk-sensitive RL in finite discounted MDPs under the recursively applied entropic risk measure, assuming that the agent is given access to a generative model of the environment. The agent’s learning performance is assessed via sample complexity defined as the total number $T$ of samples needed to learn, for input $(\varepsilon,\delta)$, either an $\varepsilon$-optimal policy (which we call policy learning), or an $\varepsilon$-close approximation (in max-norm) to the optimal Q-value (which we call Q-value learning), with probability exceeding $1-\delta$.
Specifically, we make the following contributions: We present an algorithm called Model-Based Risk-Sensitive Q-Value Iteration (MB-RS-QVI), a model-based RL algorithms for the RL problem considered, which is derived using a simple plug-in estimator. It is provably sample efficient, despite its simple design. Notably, we report sample complexity bounds on the performance of (MB-RS-QVI) under Q-value learning (Theorem 1) scaling as $\widetilde{O}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|})^{2}\Big)$ and policy learning (Theorems 2 and 3), scaling as $\widetilde{O}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{4}}(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|})^{2}\Big)$, and $\widetilde{O}\Big(\frac{S^{2}A}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\Big)$ in any discounted MDP with $S$ states, $A$ action, and discount factor $\gamma$, with $\widetilde{\mathcal{O}}$ hiding logarithmic factors. Here, $\beta$ denotes the risk parameter (see Section 2 for precise definition), where $\beta>0$ (resp. $\beta<0$) corresponds to a risk-averse (resp. risk-seeking) agent. An interesting property of these bounds is the exponential dependence on the effective horizon $\frac{1}{1-\gamma}$, which is absent in the conventional (risk-neutral) RL, wherein $\beta=0$. Another key contribution of this paper is to derive the first, to our knowledge, sample complexity lower bounds for the discounted RL under entropic risk measure. Our lower bounds establish that for Q-value learning (Theorem 4) and policy learning (Theorem 5), one needs at least $\widetilde{\Omega}\Big(\frac{(S-2)A}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\Big)$ and $\widetilde{\Omega}\Big(\frac{(S-2)(A-1)}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\Big)$ samples, respectively, to come $\varepsilon$-close to optimality. Interestingly, the derived lower bounds assert that exponential dependence on $\frac{|\beta|}{1-\gamma}$ in both sample complexities are unavoidable thus making the risk-sensitive setting fundamentally harder than the classical risk-neutral setting.
1.2 Related Work
Finite-sample guarantees for risk-neutral RL.
There is a large body of papers studying provably-sample efficient learning algorithms in discounted MDPs. These papers consider a variety of settings, including the generative setting kakade2003sample ; gheshlaghi2013minimax ; agarwal_model-based_2020 , the offline (or batch) setting rashidinejad2021bridging ; li2024settling , and the online setting strehl_analysis_2008 ; lattimore_near-optimal_2014 . In particular, in the generative setting – which we consider in this paper – some notable developments include, but not limited to kearns1998phased_Q_learning ; gheshlaghi2013minimax ; agarwal_model-based_2020 ; li2020breaking ; sidford2018variance ; wang2020randomized ; chen2020finite . Among these, gheshlaghi2013minimax ; agarwal_model-based_2020 ; sidford2018variance ; li2020breaking present algorithms attaining minimax-optimal sample complexity bounds, although some of these result do not cover the full range of $\varepsilon$. We also mention a line of work, comprising e.g. panaganti2022sample ; shi2023curious , that investigate discounted RL in the generative setting but under distributional robustness. Let us also remark that some recent works – notably zurek2024span ; zurek2025plugin – study sample complexity of average-reward MDPs in the generative setting.
Finally, we mention that some studies consider adaptive sampling in the generative setting to account for the heterogeneity across the various state-action pairs in the MDP; see, e.g., almarjani2021adaptive ; zanette2019almost . This line of works stand in contrast to the papers cited above that strive for optimizing the performance, in the worst-case sense, via uniformly sampling various state-action pairs.
@@ 1.2 Related Work
There exists a rich literature on decision making under a risk measure. We refer to sani2012risk ; maillard2013robust ; khajonchotpanya2021revised for some developments in bandits, where the performance is assessed via regret. Extensions to episodic MDPs were pursued in a recent line of work, including fei2020risk ; liang2024regret ; hu2023tighter , which establish near-optimal guarantees on regret. The two lines of work (on bandits and episodic MDPs) constitute the majority of work on risk-sensitive RL, thoroughly studying a variety of risk measures including CVaR, entropic risk, and entropic value-at-risk ahmadi2012entropic ; su2025risk .
Among the various risk measures, CVaR has arguably received a great attention; see, for instance, deng2025near ; du2023provably ; chen2024provably . For instance, du2023provably ; chen2024provably study episodic RL in the regret setting and present algorithms with sub-linear regret in tabular MDPs (du2023provably ) and under function approximation (chen2024provably ), while deng2025near investigates the sample complexity analysis in the generative setting (similar to ours), We also provide a pointer to some work which deal with a class of measures called coherent risk measures that include CVaR as a special case. In this category, we refer to, for instance, tamar2015policy , which studies policy gradient algorithms, and to lam2022risk that investigates regret minimization in the episodic setting and with function approximation. In the case of infinite-horizon setting, the entropic risk measure is mostly considered in the undiscounted (i.e., the average-reward) setting; examples include borkar2002risk ; neu2017unified ; marthe2023beyond ; marthe2025efficient . In contrast, little attention is paid to the discounted setting, which is mainly due to technical difficulties caused by discounting. The work hau2023entropic is among the few papers investigating solving discounted MDPs under the entropic risk.
@@
For $n\in\mathbb{N}$, let $n]:=\{1,\ldots,n\}$. $1_{A}$ denotes the indicator function of an event $A$. Given a set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the probability simplex over $\mathcal{X}$. We use the convention that $\|\cdot\|:=\|\cdot\|_{\infty}$ and explicitly use the subscript $\|\cdot\|_{p}$ when using $p$-norms for which $1\leq p<\infty$. We use $Z=\mathcal{S}\times\mathcal{A}$ to denote the set of all state-action pairs.
2 Background
2.1 Markov Decision Processes
We write the 6-tuple $M=(\mathcal{S},\mathcal{A},P,R,\gamma,\beta)$ as a finite, discounted infinite-horizon Markov decision process (MDP), where $\mathcal{S}=\{1,2,...,S\}$ is the finite state space of size $S:=|\mathcal{S}|$, $\mathcal{A}=\{1,2,...,A\}$ is the finite action space of size $A:=|\mathcal{A}|$, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition probability function, $R:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the deterministic reward function, $\gamma\in(0,1)$ is the discount factor, and $\beta\neq 0$ is the risk-parameter. The agent interacts with the MDP $M$ as follows. At the beginning of the process, $M$ is in some initial state $s_{0}\in\mathcal{S}$. At each time $t\geq 0$, the agent is in state $s_{t}\in\mathcal{S}$ and decides on an action $a_{t}\in\mathcal{A}$ according to some rule. The MDP generates a reward $r_{t}:=R(s_{t},a_{t})$ and a next-state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. The MDP moves to $s_{t+1}$ when the next time slot begins, and this process continues ad infinitum. This process yields a growing sequence $(s_{t},a_{t},r_{t})_{t\geq 0}$. The agent’s goal is to maximize an objective function, as a function of the reward values $(r_{t})_{t\geq 0}$ which involves the two parameters $\gamma$ and $\beta$; in addition to the discount factor $\gamma$ that makes future rewards less valuable than the present ones, the risk-parameter $\beta$ quantifies to what degree the agent seeks or avoids strategies that have more variability in the rewards obtained over time. To concretely define the objective function of the agent, we shall introduce some necessary concepts.

### v3

# Recursive Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

@@ Abstract
We study risk-sensitive reinforcement learning in finite discounted MDPs with recursive entropic risk measures (ERM), where the risk parameter $\beta\neq 0$ controls the agent’s risk attitude: $\beta>0$ for risk-averse and $\beta<0$ for risk-seeking behavior. A generative model of the MDP is assumed to be available. Our focus is on the sample complexities of learning the optimal state–action value function (value learning) and an optimal policy (policy learning) under recursive ERM. We introduce a model-based algorithm, called Model-Based ERM $Q$-Value Iteration (MB-RS-QVI), and derive PAC-type bounds on its sample complexity for both value and policy learning. Both PAC bounds scale exponentially with $|\beta|/(1-\gamma)$, where $\gamma$ is the discount factor. We also establish corresponding lower bounds for both value and policy learning, showing that exponential dependence on $|\beta|/(1-\gamma)$ is unavoidable in the worst case. The bounds are tight in the number of states and actions ($S$ and $A$), providing the first rigorous sample complexity guarantees for recursive ERM across both risk-averse and risk-seeking regimes.
@@ 1 Introduction
In reinforcement learning (RL), the aim of the agent is to conventionally maximize the expected return, which is defined in terms of a (possibly discounted) sum of rewards [71]. The environment is typically modeled via the Markov Decision Process (MDP) framework [60], wherein efficient computation of an optimal policy, thanks to optimal Bellman equations, renders possible. However, as a risk-neutral objective, the expected return is inadequate to capture the true needs of many high-stake applications arising in, e.g., medical treatment [24], finance [65, 10], operations research [20], and transportation [39]. In these domains, decision making must take into account the variability of returns, and risks thereof. To address this limitation, one may opt to maximize a risk measure of the return distribution. Alternatively, one may model the entire distribution of return, as in distributional RL [9], which has received significant attention over the last decade. In this paper, we focus on the former.
@@
Within the first approach, the risk is elegantly quantified via concave risk measures, which lead to well-defined optimization problems. Notable risk measures include mean-variance [46], value-at-risk (VaR) [21], Conditional VaR (CVaR) [66], entropic risk [33], and entropic VaR (EVaR) [2], all of which have been applied to a wide-range of scenarios. Among these, CVaR has become particularly popular for modeling risk-sensitivity in MDPs [19, 11, 15, 7], mainly due to a delicate control it offers for the undesirable tail of return distribution. Despite its popularity and appealing interpretation, learning in MDPs with CVaR-based objectives may pose technical challenges [7]. ERM, as another popular notion, has long been considered for risk-sensitive control in MDPs and RL [33, 12, 32, 34, 25]. However, much of the existing literature focuses on undiscounted settings, despite the prevalence of discounted MDPs; see, e.g., [8, 32, 55] for notable exceptions.
In risk-sensitive RL with a specified risk measure, objectives can be formulated in two fundamentally different ways, depending on how the risk functional is applied to the reward sequence $(r_{t})_{t\geq 0}$. The first approach, referred to as non-recursive (also called non-iterated or static), consists in directly applying the risk functional to the total return (e.g., $\sum_{t=0}^{\infty}\gamma^{t}r_{t}$ in the discounted case) [14, 12, 31]. The second, termed recursive (also called iterated, nested, or dynamic), applies the risk functional at every step $t$ to the reward-to-go [4, 5, 22]; see Section 3 for details. In other words, the non-recursive formulation captures trajectory-level risk, whereas the recursive formulation deals with risk locally at each step. The two formulations are qualitatively different and should be viewed as orthogonal modeling choices. While trajectory-level risk is often easier to interpret, it may lead to policies that allow the agent to visit high-risk states, even though the risk of the entire trajectory is still controlled, which might be unacceptable in many safety-critical applications. In contrast, the recursive formulation may lead to more cautious behavior by discouraging entry into high-risk states at every step, which can be either desirable or overly conservative depending on the application [22]. From a theoretical perspective, another key distinction is that non-recursive formulations do not generally admit Bellman-type optimality equations and may lead to time-inconsistent optimal policies (see [36]), whereas recursive formulations preserve Bellman-type optimality structures. Motivated by these considerations, we study risk-sensitive discounted RL with objectives defined via the recursive ERM.
1.1 Main Contributions and Paper Organization
@@
We consider risk-sensitive RL in tabular discounted MDPs under recursive ERM, encompassing both the risk-averse and risk-seeking regimes. We assume that the agent has access to a generative model of the MDP, namely, a simulator that generates samples from the true MDP for arbitrary state-action pairs. Learning performance is assessed in terms of sample complexity, defined as the total number $T$ of samples required, for given $(\varepsilon,\delta)$, to obtain either an $\varepsilon$-optimal policy (policy learning) or an $\varepsilon$-close approximation of the optimal Q-value in the max-norm (value learning), with probability exceeding $1-\delta$.
We make the following contributions. We develop a model-based algorithm, called Model-Based ERM Q-Value Iteration (MB-RS-QVI), and establish PAC-type bounds on its sample complexity for both value learning and policy learning. MB-RS-QVI is based on a plug-in estimation of the transition kernel combined with a Q-value iteration scheme adapted to recursive ERM objectives. This QVI structure is inspired by the value iteration method of [5], which considers the risk-averse planning setting with known dynamics. We show that its sample complexity for value learning (Theorem 1) and policy learning (Theorem 2) scales as [^3] Throughout the paper, $\widetilde{\mathcal{O}}$ and $\widetilde{\Omega}$ suppress logarithmic factors in the relevant problem parameters.
$\widetilde{O}\bigg(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}|\beta|^{2}}e^{2|\beta|/(1-\gamma)}\bigg)\qquad\hbox{and}\qquad\widetilde{O}\bigg(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}|\beta|^{2}}\min\Big\{S,\frac{1}{(1-\gamma)^{2}}\Big\}e^{2|\beta|/(1-\gamma)}\bigg)\,$,
respectively. These bounds hold for any discounted MDP with $S$ states, $A$ actions, discount factor $\gamma$, and risk parameter $\beta$, with $\beta>0$ (respectively, $\beta<0$) corresponding to a risk-averse (respectively, risk-seeking) agent; see Section 3 for a precise definition. Moreover, the bounds are valid over the entire $\varepsilon$-range, namely $\varepsilon\in(0,\frac{1}{1-\gamma}]$. A notable feature of these results is the exponential dependence on the effective horizon $1/(1-\gamma)$, which is absent in the conventional risk-neutral setting, wherein $\beta=0$.
We further establish worst-case lower bounds on the sample complexity of recursive ERM. Specifically, we show that for value learning (Theorem 3) and policy learning (Theorem 4), at least
$\widetilde{\Omega}\bigg(\frac{SA}{\varepsilon^{2}|\beta|^{2}}e^{|\beta|/(1-\gamma)}\bigg)$
samples are required to achieve $\varepsilon$-optimality. These lower bounds demonstrate that the exponential dependence on $|\beta|/(1-\gamma)$ in sample complexity upper bounds is unavoidable, thereby establishing that learning under recursive ERM is fundamentally more challenging than in the risk-neutral case. To the best of our knowledge, these results constitute the first upper and lower bounds on the sample complexity of recursive ERM in discounted MDPs. A summary of our results is provided in Table 1.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 introduces the necessary background and formal problem setup. Section 4 presents the MB-RS-QVI algorithm, while Section 5 reports its sample complexity guarantees, with proofs deferred to Section 6. Lower bounds are presented in Section 7. Section 8 presents numerical results to demonstrate the performance of MB-RS-QVI. Finally, Section 9 concludes with a discussion and directions for future research. Additional background on risk measures, along with omitted proofs, is provided in the appendix.
Problem Upper Bound Lower Bound
ERM (value learning) $\widetilde{\mathcal{O}}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}|\beta|^{2}}e^{2|\beta|/(1-\gamma)}\Big)$    [Theorem 1] $\widetilde{\Omega}\Big(\frac{SA}{\varepsilon^{2}|\beta|^{2}}e^{|\beta|/(1-\gamma)}\Big)$    [Theorem 3]
ERM (policy learning) $\widetilde{\mathcal{O}}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}|\beta|^{2}}\min\big\{S,\frac{1}{(1-\gamma)^{2}}\big\}e^{2|\beta|/(1-\gamma)}\Big)$    [Theorem 2] $\widetilde{\Omega}\Big(\frac{SA}{\varepsilon^{2}|\beta|^{2}}e^{|\beta|/(1-\gamma)}\Big)$   [Theorem 4]
Risk-neutral (value learning) $\widetilde{\mathcal{O}}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{3}}\Big)$ [30] $\widetilde{\Omega}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{3}}\Big)$ [30]
Risk-neutral (policy learning) $\widetilde{\mathcal{O}}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{3}}\Big)$ [30, 1, 49, 37] $\widetilde{\Omega}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{3}}\Big)$ [30]
Table 1: Summary of upper and lower bounds presented in this paper. $\beta$ denotes the risk parameter, where $\beta>0$ (respectively, $\beta<0$) corresponds to a risk-averse (respectively, risk-seeking) agent.
2 Related Work
Risk-neutral discounted RL.
There is a large body of papers on provably-sample efficient learning algorithms in tabular discounted MDPs, encompassing a variety of settings such as the generative setting [38, 30, 1, 67, 49, 37], the offline (or batch) setting [61, 47], and the online setting [70, 45]. In our overview of risk-neutral work, we restrict attention to the generative setting —which is the setting we consider— with the aim of collecting most notable developments and key results. In this line, [40] reports the earliest known sample complexity bounds, which is achieved by a model-free method. Azar et al. [30] substantially improve upon this by showing that a simple model-based method attains optimal sample complexity bounds scaling as $\widetilde{\cal O}\big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{3}}\big)$ for both value learning and policy learning, albeit for substantially limited $\varepsilon$-ranges. It also establishes a lower bound of $\widetilde{\Omega}\big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{3}}\big)$ for value learning. Further algorithms and results are reported in more recent subsequent work, notably including [67, 74, 48, 37]. Among these, [67, 74, 37] present model-free methods, with [67, 37] presenting minimax-optimal bounds, although valid for restricted $\varepsilon$-ranges. Under model-based methods, minimax-optimal bounds, beyond [30], are reported in [1, 48]. In particular, [1] uses empirical MDP combined with a black-box planner, and reports a minimax-optimal bound for $\varepsilon\in\big(0,\frac{1}{\sqrt{1-\gamma}})$, thus expanding that in [30]. More recently, [48] establishes minimax-optimal bounds for the entire $\varepsilon$-range, which are achieved by model-based methods built via the empirical MDP but with reward perturbations or conservative planning. It is worth emphasizing that existing optimal sample complexities rely on techniques that crucially exploit the additivity of the return with respect to rewards; this structural property generally fails for risk-sensitive objectives, and the corresponding techniques do not carry over.
We remark that the abovementioned works look at the learning performance in a worst-case scenario, which yield sample complexity bounds that hold for a model class. This is typically done via uniformly sampling various state-action pairs. In contrast, some studies (e.g., [3, 76]) consider adaptive sampling to account for the heterogeneity across state-action space of the MDP, typically resulting in instance-dependent bounds.
@@ 2 Related Work
There exists a substantial literature on decision making under a risk measure in bandit and RL settings. In bandits, risk-sensitive objectives are typically studied through regret minimization; see, e.g., [64, 51, 42]. Extensions to MDPs introduce substantially richer structural and algorithmic challenges, which are the focus of this work. The literature on RL under risk measures may be broadly categorized by the type of risk measure studied. Representative examples include CVaR [22, 23, 16, 44], ERM [12, 56, 53, 52, 32], mean-variance risk [69, 35, 43], and EVaR [57, 29]. Among these, CVaR has been the most extensively studied. Under non-recursive CVaR, [23] and [16] investigate online episodic RL in the regret setting for tabular MDPs and MDPs with function approximation, respectively, while reward-free RL is studied in [58]. Under recursive CVaR, [22] analyzes sample complexity in the generative setting; however, its analysis relies on structural properties specific to CVaR and does not extend to ERM.
ERM has also been widely studied, beginning with early work such as [33] and followed by a rich literature across diverse settings [12, 13, 14, 63, 25, 27, 26, 53, 52, 56, 32]. Under non-recursive ERM, most work focuses on the average-reward or episodic settings, with planning studied in [33, 12, 52] and learning in [14, 56, 53]. The discounted setting has received comparatively little attention, largely due to technical challenges introduced by discounting; notable exceptions include planning results in [8, 32] and learning results in [55], which modify the problem formulation to address time inconsistency. Under recursive ERM, recent works such as [25, 26, 27, 34, 50] study online episodic RL in the regret setting. To the best of our knowledge, existing work on discounted MDPs under recursive ERM is limited to planning; a notable example is [5], which provides a thorough theoretical treatment but does not propose learning algorithms. The analysis in [5] (and the works in the known-model setting) is limited to the case of known models, where the problem does not involve statistical estimation. As a result, sample complexity analyses (under policy or value learning), which aim to characterize statistical hardness, are not relevant in this setting.
Some work in risk-sensitive RL and control studies broader classes of risk measures. Two notable classes studied in this context are coherent risk measures and optimized certainty equivalent (OCE) measures, both of which include important special cases such as mean–variance and CVaR. While ERM is not coherent, it belongs to the OCE class for $\beta>0$; a brief overview of risk measures is provided in Appendix A. Results for coherent risks [59, 72, 44, 77] do not apply to ERM, and existing results for OCE risks [73, 75, 62] do not address provably sample-efficient learning under recursive ERM in discounted MDPs. In particular, [62] considers offline RL in discounted MDPs under recursive OCE but does not provide sample-complexity guarantees. We also note that a connection between MDPs with recursive coherent risks and distributionally robust MDPs has been established in [5]. Finally, we note that approaches such as safe RL and constrained MDPs [18, 17] incorporate risk awareness into policies via constraints, but without altering the definition of return; they are therefore generally regarded as orthogonal to the present setting.
3 Background
@@
For $n\in\mathbb{N}$, let $n]:=\{1,\ldots,n\}$. $\mathbbmss{1}_{A}$ denotes the indicator function of an event $A$. Given a set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the probability simplex over $\mathcal{X}$. We use the convention that $\|\cdot\|:=\|\cdot\|_{\infty}$ and explicitly use the subscript $\|\cdot\|_{p}$ when using $p$-norms for $1\leq p<\infty$. The notation $L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ denotes the space of essentially bounded random variables on the probability space $(\Omega,\mathcal{F},\mathbb{P})$.